Dfuentes642:

The temperature at a mountain base camp was −4 degrees Celsius on Saturday. Sunday morning, the temperature was 2 degrees Celsius lower than it was on Saturday. By Sunday evening, the temperature was 1 degree Celsius lower than it was that morning. What was the temperature at the base camp Sunday evening? −9 degrees Celsius −8 degrees Celsius −7 degrees Celsius −5 degrees Celsius

Gdeinward:

solve this equation: (-4 - 2) - 1=

Dfuentes642:

ok

Dfuentes642:

give me a minute or two.

Gdeinward:

alright.

Dfuentes642:

-7

Gdeinward:

that is correct. GOOD JOB!!!

Dfuentes642:

ok whats next

Gdeinward:

that is your answer

Dfuentes642:

or is that straight up the answer

Dfuentes642:

A seagull can dive 45 meters below the surface of the water to reach a fish. It dives 11% of the total depth every 1 second. What is the depth of the seagull after 2 seconds? –2.475 meters –4.95 meters –9.9 meters –7.425 meters

Gdeinward:

All I did was take each amount subtracted from the origional, put it in equation form and then viola

Dfuentes642:

this is next question

Gdeinward:

Do you know how to calculate percentages?

Dfuentes642:

NO

Dfuentes642:

no

Dfuentes642:

no

Gdeinward:

Simpy take the percentage, put it in decimal form, and multiply by the total: 11 = 0.11, and 45 times 0.11

Dfuentes642:

so 45 timea 0.11

Dfuentes642:

that iws what i do cause im confused

Gdeinward:

which is 4.95, then you subtract that from the 45

Dfuentes642:

what?

Dfuentes642:

im not that smart

Gdeinward:

the seagull dives 45 meters, and 11 percent of the total depth each second.

Dfuentes642:

what do i do

Dfuentes642:

multiply?

Gdeinward:

so to find 11 percent of 45, you multiply 45 by 0.11 to get 4.95

Gdeinward:

so in 1 second he dives 4.95 feet.

Dfuentes642:

so thats the answer?

Dfuentes642:

Use long division to determine the decimal equivalent of 8 over 9. 0 point 3 bar 0 point 4 bar 0 point 5 bar 0 point 8 bar

Dfuentes642:

next questioon

Gdeinward:

then to figure out the next total depth, we subtract 4.95 from 45, because we want to know how deep he is after 2 seconds

Gdeinward:

we are not done with the last one.

Dfuentes642:

oh

Gdeinward:

next simply follow the steps again. 40.05 is what you get when you subtract 4.95 from 45.

Dfuentes642:

ok but that is not the options

Gdeinward:

so we must again find 11 percent of 40.05. Do you know how?

Dfuentes642:

no

Dfuentes642:

hello

Gdeinward:

If you saw what we did before, to find a percentage of a number, we turn the percent into a decimal. Then multiply the number by that decimal

Gdeinward:

so 40.05 times 0.11.

Dfuentes642:

ok

Dfuentes642:

4.445

Gdeinward:

so what is 40.05 minus 4.445?

Dfuentes642:

35.605

Gdeinward:

okay so we know what his depth is now, but they want to know how far from sea level, so we subtract 35.605 from 45.

Gdeinward:

which is?

Dfuentes642:

-9.395

Gdeinward:

your answer is C.

Dfuentes642:

ok

Dfuentes642:

Simplify negative 3 over 5 ÷ 7 over 6. negative 7 over 10 negative 18 over 35 18 over 35 7 over 10

Dfuentes642:

please hurry becaus ei have family business to attend

Dfuentes642:

we are almost done

Gdeinward:

so -3/5 / 7/6. Do you know how to go from fraction to decimal form?

Dfuentes642:

no

Dfuentes642:

hello

Dfuentes642:

hello

Gdeinward:

you must divide the top number by the bottom number

Dfuentes642:

ok

Gdeinward:

so -3/5 = -0.6

Gdeinward:

and 7/6 = 1.16

Gdeinward:

so -0.6 divided by 1.16 = -0.51724137931

Dfuentes642:

i still dont understand

Gdeinward:

you need to change all the fractions in the answer choices to decimals, and choose the answer that has the same decima that we just found.

Dfuentes642:

they arent decimals though

Gdeinward:

yeah but you turn the fractions into decimals like we did before.

Dfuentes642:

which is how

Gdeinward:

the answer is B.

Dfuentes642:

ok

Dfuentes642:

thanks

Gdeinward:

divide the top nuber by the bottom

Dfuentes642:

Convert 7 over 8 to a decimal. 0.375 0.875 0.925 0.965

Gdeinward:

Hey hey, like we did before, divide the top number by the bottom.

Dfuentes642:

ok

Dfuentes642:

0.875

Gdeinward:

so the answer is B.

Dfuentes642:

ok

Dfuentes642:

Convert 1 over 8 to a decimal and tell whether it terminates or repeats. 0.125, a number that terminates 1.8, a number that terminates 1.8 with a repeating bar over 8., a number that repeats 0.875 with a repeating bar over 875., a number that repeats

Gdeinward:

so turn it into a decimal like we did before. Terminating numbers dont have decimals with the same repeating number, and non terminating decimas do.

Dfuentes642:

can u please clearly explain to me again

Gdeinward:

to turn a fraction into a decimal, divide the top number by the bottom.

Gdeinward:

do that and tell me what you get.

Dfuentes642:

0.125

Gdeinward:

Tell what the decimal is doing. Does it have any repeating numbers?

Dfuentes642:

no

Gdeinward:

then it is a terminating decimal number.

Gdeinward:

A non terminating decimal is a decimal that is repeating.

Gdeinward:

so the answer is A

Dfuentes642:

ok

Dfuentes642:

Jacob is showing his steps to solve the expression −27.8 ÷ 1 over 7. In which step did Jacob make an error? Step 1: −27.8 ÷ 1 over 7 Step 2: −27.8 ⋅ 7 Step 3: −194.6 Step 1 Step 2 Step 3 There is no error.

Gdeinward:

Do you have any clue?

Dfuentes642:

no

Gdeinward:

there is no error.

Dfuentes642:

ok

Dfuentes642:

Matt is making cookies. He adds 3 over 4 cup of sugar to a bowl. He then removes two 1 over 8 cups of sugar. Which of the following best explains the amount of sugar left in the bowl? 3 over 8, because 3 over 4 minus 2 times 1 over 8 equals 3 over 4 minus 2 over 8 3 over 8, because 3 over 4 minus 2 times 1 over 8 equals 3 over 4 minus 3 over 8 4 over 8, because 3 over 4 minus 2 times 1 over 8 equals 3 over 4 minus 3 over 8 4 over 8, because 3 over 4 minus 2 times 1 over 8 equals 3 over 4 minus 2 over 8 You must check the box below prior to submitting your exam! Check this box to indicate you are ready to submit your exam Submit Exam Explorer Toolbox Workload My Folders Current course: Instructors monitor ALL areas of a student's account Student e-mail accounts are to be used for FLVS course-related email only and not for general introductions or spamming of people in your address book. Please remember to click the Logoff link when you have completed your work in the course. FDK391.03 Previous QuestionNext Question

Gdeinward:

what is 3/4 - 2/8 ?

Dfuentes642:

i dont kn ow

Gdeinward:

you must utelize gcf turn 3/4 into 6/8: 3/4. 3 times 2 = 6 4 times 2 = 6 3/4 = 6/8 6/8 - 2/8 = 4/8

Dfuentes642:

so which one

Gdeinward:

the answer is D.

Dfuentes642:

thank you

Dfuentes642:

i will tell you what i get

Gdeinward:

Your Welcome.

Gdeinward:

sounds good

Dfuentes642:

83.33 percent

Dfuentes642:

a B

Gdeinward:

For future reference, we are not allowed to complete tests for you. Did you do any on your own, or did I answer every question?

Gdeinward:

Thanks Evo, but thats not necessary.

Gdeinward:

d